### 08 Mar 2024 v1

@@ Abstract
•  Data Construction: We strive to ensure our data is diverse, scalable and extensively covers real-world scenarios including web screenshots, PDFs, OCR, charts, and knowledge-based content (expert knowledge, textbooks), aiming for a comprehensive representation of practical contexts. Further, we create a use case taxonomy from real user scenarios and construct an instruction tuning dataset accordingly. The fine-tuning with this dataset substantially improves the model’s user experience in practical applications.
@@ 1 Introduction
  • While some models (Sun et al., 2023; Wang et al., 2023b; 01-ai, 2024; Lin et al., 2023) have begun to exploit pretraining, they often overlook the preservation of language skills. Often, there is a degradation of language capability after prolonged multimodal training. Since we aim for a generalist that possesses strong capabilities in both modalities, there should be a training strategy that well preserves the language capability when developing the new modality ability.
@@
When iterating on our model, We conduct experiments on a small scale before scaling to a larger model size. However, a smaller model, e.g.,, 1B model, cannot demonstrate reasonable performance on benchmarks (Schaeffer et al., 2024) and faithfully reflect the model’s performance. We adopt two approaches to address this. First, we modify the evaluation protocol from multi-choice to compare the perplexity of options. Also, to prevent the instruction following ability from becoming the bottleneck, we mix a small proportion of instruction tuning data during the pretraining phase. In this way, we can achieve reasonable performance using the 1B model and more accurately measure the impact of each iteration during the experiment.
@@ 2.1 Vision-Language pretraining Data
Interleaved image-text MMC4 (Zhu et al., 2024) 15.2%
@@
Document OCR arXiv rendered markdown
@@
Interleaved image-text data enable the models to have a better capability for in-context learning of multi-modality inputs, and we utilize three public datasets MMC4 (Zhu et al., 2024), Wiki (Burns et al., 2023) and Wikihow (Yang et al., 2021).
@@
Table and chart data enable the models to learn the capability for general table and chart image understanding. It encompasses a diverse range of public data sources, including Chart2text (Kantharaj et al., 2022), Geo170K (Gao et al., 2023), Unichart (Masry et al., 2023), Ureader (Ye et al., 2023), M-paper (mPLUG, 2024), ScienceQA (Lu et al., 2022b), ScreenQA (Hsiao et al., 2022), SciGraphQA-295K (Li and Tajbakhsh, 2023), Paper2figure100k (Rodriguez et al., 2023), Widget Captioning (Li et al., 2020), Screen2words (Wang et al., 2021), and Refexp (Mao et al., 2016).
@@ 3.1 Architecture
To address these limitations, we additionally utilize a vision-only encoder based on the SAM-B (Kirillov et al., 2023), a pre-trained ViTDet (Li et al., 2022) image encoder to process low-level features, which accepts high-resolution 1024 x 1024 image inputs. In addition to the SAM-B encoder, we retain the SigLIP-L vision encoder with low-resolution 384 x 384 image inputs. Consequently, our hybrid vision encoder combines the SAM-B and SigLIP-L encoders, efficiently encoding high-resolution 1024 x 1024 images while preserving both semantic and detailed information. Specifically, a high-resolution SAM-B vision encoder first resizes the image into 1024 x 1024 and results in a 64 x 64 x 256 feature map. Then we will interpolate it into 96 x 96 x 256 and employ two convolutional layers with a stride of 2, producing a feature map of 24 x 24 x 1024, and finally reshape it to 576 x 1024. The low-resolution SigLIP-L vision encoder first resizes the image into 384 x 384, encodes it into a 24 x 24 x 1024 feature map, and finally reshapes it into 576 x 1024.
@@ 3.2.2 Stage 2: Joint Vision-Language pretraining
Figure 4: Comparative performance results on different modality fusion ratio. An excessively large proportion of multimodal data (multimodal:language=100%:0%) leads to significant forgetting of language capabilities in LLMs. A suitable ratio (multimodal:language=70%:30%) can effectively mitigate the issue of language forgetting while simultaneously enhancing the model’s multimodal abilities.
@@ 3.3 Hyperparameters and Infrastructures
The detailed hyperparameters of all stages are illustrated in Table 4. We train and evaluate our DeepSeek-VL with HAI-LLM (High-flyer, 2023), a lightweight and efficient distributed training framework. Since we use visual encoders to convert images into embedding vectors and then treat image embeddings and text embeddings uniformly, we can easily adapt pipeline parallelism to VL model training: all we need to do is to view visual encoders and text embedding as a single module and take it as the first layer of the resulting model. This very first layer has a complicated model structure and precludes standard tensor parallelism technique, but luckily it requires relatively small computation compared to upper standard transformer blocks. We therefore simply recompute the visual encoder forward pass in all tensor parallel ranks. The existence of visual encoders also leads to non-uniform execution time across model layers, so we re-divide model layers between pipeline parallelism ranks to achieve better load balance and throughput. The upper layers of DeepSeek-VL are exactly the same as those in DeepSeek-LLM. With such minor modification, we can now perform canonical 3D parallelism techniques as in Megatron (Shoeybi et al., 2019; Narayanan et al., 2021; Korthikanti et al., 2023) and overlap computation and communication as in DeepSeek-LLM (DeepSeek-AI, 2024). DeepSeek-VL-7B consumed 5 days on a cluster of 64 Nvidia A100 processing units, each comprising 8 cards, while DeepSeek-VL-1B consumed 7 days on a setup involving 16 Nvidia A100 processing units.
DeepSeek-VL 1B DeepSeek-VL 7B
@@ 4.3 Human Evaluation
Figure 6: Human evaluation results on InternLM2-VL, CogVLM, DeepSeek-VL and GPT-4V.
We compare our DeepSeek-VL 7B with InternLM2-VL, CogVLM and GPT-4V as shown in Figure 6 (and we also provide visualization results in Appendix A). GPT-4V demonstrates exceptional performance across most dimensions. All open-source models are still far behind GPT-4V in logical reasoning, highlighting the necessity of scaling up the size of Large Language Models (LLMs). DeepSeek-VL 7B achieves better results in overall performance, reaching outcomes close to GPT-4V in Recognition, Conversion, and Commonsense Reasoning.
@@ 4.4 Ablation Study
Figure 7: Comparative analysis of modality warmup on language (Pile-test) and multimodal (MMBench and MMBench__\__CN) benchmarks demonstrates that modality grouping consistently surpasses the non-grouped modality approach in language tasks, while simultaneously preserving performance on multimodal tasks (Multimodal:Language=60%:40%).
@@
Figure 8: Comparative performance results on language (Pile-test) and multimodal (MMBench and MMBench_CN) benchmarks for modality warmup. Modality warmup consistently matches or surpasses the performance of approaches without modality warmup across all evaluated tasks (Multimodal:Language=60%:40%).

### 11 Mar 2024 v2

@@ Abstract
•  Data Construction: We strive to ensure our data is diverse, scalable and extensively covers real-world scenarios including web screenshots, PDFs, OCR, charts, and knowledge-based content (expert knowledge, textbooks), aiming for a comprehensive representation of practical contexts. Further, we create a use case taxonomy from real user scenarios and construct an instruction-tuning dataset accordingly. The fine-tuning with this dataset substantially improves the model’s user experience in practical applications.
@@ 1 Introduction
  • While some models (Sun et al., 2023; Wang et al., 2023b; 01-ai, 2024; Lin et al., 2023a) have begun to exploit pretraining, they often overlook the preservation of language skills. Often, there is a degradation of language capability after prolonged multimodal training. Since we aim for a generalist that possesses strong capabilities in both modalities, there should be a training strategy that well preserves the language capability when developing the new modality ability.
@@
When iterating on our model, We conduct experiments on a small scale before scaling to a larger model size. However, a smaller model, e.g., 1B model, cannot demonstrate reasonable performance on benchmarks (Schaeffer et al., 2024) and faithfully reflect the model’s performance. We adopt two approaches to address this. First, we modify the evaluation protocol from multi-choice to compare the perplexity of options. Also, to prevent the instruction following ability from becoming the bottleneck, we mix a small proportion of instruction tuning data during the pretraining phase. In this way, we can achieve reasonable performance using the 1B model and more accurately measure the impact of each iteration during the experiment.
@@ 2.1 Vision-Language pretraining Data
Interleaved image-text MMC4 (Zhu et al., 2024) 13.1%
@@
Document OCR arXiv rendered markdown (Blecher et al., 2023) 2.1%
@@
Interleaved image-text data enable the models to have a better capability for in-context learning of multi-modality inputs, and we utilize three public datasets MMC4 (Zhu et al., 2024), Wiki (Burns et al., 2023), Wikihow (Yang et al., 2021) and Epub textbooks.
@@
Table and chart data enable the models to learn the capability for general table and chart image understanding. It encompasses a diverse range of public data sources, including Chart2text (Kantharaj et al., 2022), Geo170K (Gao et al., 2023), Unichart (Masry et al., 2023), Ureader (Ye et al., 2023), M-paper (Hu et al., 2023), ScienceQA (Lu et al., 2022b), ScreenQA (Hsiao et al., 2022), SciGraphQA-295K (Li and Tajbakhsh, 2023), Paper2figure100k (Rodriguez et al., 2023), Widget Captioning (Li et al., 2020), Screen2words (Wang et al., 2021), and Refexp (Mao et al., 2016).
@@ 3.1 Architecture
To address these limitations, recent researches (Wei et al., 2023; Tong et al., 2024; Lin et al., 2023b) have advocated for the integration of additional vision-only self-supervised encoders, to enhance the visual grounding capabilities of multi-modality models. Building upon previous motivations, we additionally utilize a vision-only encoder based on the SAM-B (Kirillov et al., 2023), a pre-trained ViTDet (Li et al., 2022) image encoder to process low-level features, which accepts high-resolution 1024 x 1024 image inputs. In addition to the SAM-B encoder, we retain the SigLIP-L vision encoder with low-resolution 384 x 384 image inputs. Consequently, our hybrid vision encoder combines the SAM-B and SigLIP-L encoders, efficiently encoding high-resolution 1024 x 1024 images while preserving both semantic and detailed information. Specifically, a high-resolution SAM-B vision encoder first resizes the image into 1024 x 1024 and results in a 64 x 64 x 256 feature map.
@@ 3.2.2 Stage 2: Joint Vision-Language pretraining
Figure 4: Comparative performance results on different modality fusion ratio on training stage 2. An excessively large proportion of multimodal data (multimodal:language=100%:0%) leads to significant forgetting of language capabilities in LLMs. A suitable ratio (multimodal:language=70%:30%) can effectively mitigate the issue of language forgetting while simultaneously enhancing the model’s multimodal abilities.
@@ 3.3 Hyperparameters and Infrastructures
The detailed hyperparameters of all stages are illustrated in Table 4. We train and evaluate our DeepSeek-VL with HAI-LLM (High-flyer, 2023), a lightweight and efficient distributed training framework. Since we use visual encoders to convert images into embedding vectors and then treat image embeddings and text embeddings uniformly, we can easily adapt pipeline parallelism to VL model training: all we need to do is to view visual encoders and text embedding as a single module and take it as the first layer of the resulting model. This very first layer has a complicated model structure and precludes standard tensor parallelism technique, but luckily it requires relatively small computation compared to upper standard transformer blocks. We therefore simply recompute the visual encoder forward pass in all tensor parallel ranks. The existence of visual encoders also leads to non-uniform execution time across model layers, so we re-divide model layers between pipeline parallelism ranks to achieve better load balance and throughput. The upper layers of DeepSeek-VL are exactly the same as those in DeepSeek-LLM. With such minor modification, we can now perform canonical 3D parallelism techniques as in Megatron (Shoeybi et al., 2019; Narayanan et al., 2021; Korthikanti et al., 2023) and overlap computation and communication as in DeepSeek-LLM (DeepSeek-AI, 2024). DeepSeek-VL-7B consumed 5 days on a cluster of 64 nodes, each comprising 8 Nvidia A100 GPUs, while DeepSeek-VL-1B consumed 7 days on a setup involving 16 nodes.
DeepSeek-VL 1B DeepSeek-VL-7B
@@ 4.3 Human Evaluation
Figure 6: Human evaluation results on InternLM-XComposer2-VL (Dong et al., 2024), CogVLM (Wang et al., 2023b), DeepSeek-VL and GPT-4V (OpenAI, 2023b).
We compare our DeepSeek-VL-7B with InternLM-XComposer2-VL, CogVLM and GPT-4V as shown in Figure 6 (and we also provide visualization results in Appendix A). GPT-4V demonstrates exceptional performance across most dimensions. All open-source models are still far behind GPT-4V in logical reasoning, highlighting the necessity of scaling up the size of Large Language Models (LLMs). DeepSeek-VL-7B achieves better results in overall performance, reaching outcomes close to GPT-4V in Recognition, Conversion, and Commonsense Reasoning.
Figure 7: GPT-4V-based Evaluation Results of DeepSeek-VL vs. Other Models: The chart depicts results from a GPT-4V-based assessment across 99 test samples, demonstrating DeepSeek-VL’s favorable outcomes against both open-source and proprietary models.
In addition, we conduct a comparative assessment using GPT-4V to evaluate the performance of DeepSeek-VL against other models across a set of 99 test samples designed for human evaluation. Following (Zheng et al., 2024), we show GPT-4V the question and the answers from two different models and ask GPT-4V to determine which one is better or declare a tie. The results indicate a preference for DeepSeek-VL’s responses in the majority of cases, as GPT-4V tends to rate the quality of DeepSeek-VL’s answers more favorably. As illustrated in Figure 7, DeepSeek-VL is judged to be superior in over 60% of instances when compared to open-source multimodal models, including Fuyu-8B, CogVLM-17B, and InternLM-XComposer2-VL. Moreover, in comparison with other proprietary models, such as GPT-4V itself, DeepSeek-VL demonstrates comparably exceptional performance.
@@ 4.4 Ablation Study
Figure 8: Comparative analysis of modality warmup on language (Pile-test) and multimodal (MMBench and MMBench__\__CN) benchmarks demonstrates that modality grouping consistently surpasses the non-grouped modality approach in language tasks, while simultaneously preserving performance on multimodal tasks on training stage 2 (Multimodal:Language=60%:40%).
@@
Figure 9: Comparative performance results on language (Pile-test) and multimodal (MMBench and MMBench_CN) benchmarks for modality warmup. Modality warmup consistently matches or surpasses the performance of approaches without modality warmup across all evaluated tasks on training stage 2 (Multimodal:Language=60%:40%).